Suppose the probability of selling a car today is 0.36. Find the odds against selling a car?
Answer by stanbon(48516)

Suppose the probability of selling a car today is 0.36. Find the odds against selling a car?
-------------------------
Definition: "odds against A" = P(not A)/P(A)
-------------------------
P(not selling a car today) = 0.64
---
Odds against selling a car today = 0.64/0.36 = 16:9
=========================================================
Cheers,
Stan H.